OpenStudy (anonymous):

the integral 3/x^2-1 between 1 and 2, again i am having some a problems with it, i know the answer is 1/2, but from what my adding and subtracting i get, i get a different answer, and you have to use the first fundamental theorem of calculus please help

OpenStudy (turingtest):

can you show your working so far?

OpenStudy (anonymous):

\[\int\limits_{1}^{2}\frac{ 3 }{x^{2}}-1=\int\limits_{1}^{2} (3x^{-1}-x)\]

OpenStudy (anonymous):

\[\int \frac{dx}{x^2}=-\frac{1}{x}+C\]